Status Report

NOAA GOES SATOPS Morning Report 5 Feb 2001

By SpaceRef Editor
February 5, 2001
Filed under ,


GOES-8 Rapid Scan Operations approved 1326z – 2126z on February 2. Requested
by Marquette Michigan for Lake Effect Snow.

Senior Duty Meteorologist declared Critical Weather days for 36 hours
from 0000z February 5 to 1200z February 6.



NOAA-15 rev 14182 / F at 1933z on February 3: G5B lost due to ground
recorder crash at CDA. 100% of G5B recovered on rev 14185.

NOAA-16 revs 1903, 1908 on February 3 and rev 1916 on February 4: Recorder
5B did not appear to be at beginning of tape at the conclusion of the scheduled
GAC playbacks. It was positioned manually.

NOAA-16 rev 01921 / F at 1529z on February 4: According to the data
quality analysis, only 93.6% of the expected data was received, and the
CEMSCS monitor indicated that at least 6 minutes of data was missing.

NOAA-16 rev 01922 / W at 1655z on February 4: Replayed G5B from previous
pass to determine if all expected data was actually recorded on the recorder
and could be recovered.   92.8% of the expected data was received
and the same 6 minutes of data missing on the previous pass, which should
have been  recorded near beginning of tape, was missing.  Additionally,
crews noted that frame count stopped increasing at least 10 seconds prior
to the recorder reaching beginning of tape.  Motor current was intermittently
flagging out of limits.  Engineering notified.

NOAA-16 Digital Tape Recorder (DTR) 5B anomaly developed and was investigated
in depth on February 4. Current limit violations resulted in crews notifying
engineering and they began troubleshooting. Approximately the first 10
seconds of tape were getting “lost” from each recorder 5B dataset. Test
ingests supported by CEMSCS revealed no data was recorded.  Scheduling
support was on-site and has removed recorder 5B from schedule as of the
February 2 load – it will remain out of the recording sequence until the
problem is resolved.  Engineering will investigate further and Operations
will continue to monitor.



No change in the status of the DMSP spacecraft.

DMSP operations were nominal over the past 72 hours.



No change in the status of the GOES spacecraft.

At 2006z on February 2: GOES-8 image frame break due to loss of signal

At 0709z on February 3: GOES-8 Southern hemisphere canceled for dual
cord operations.

At 1346z on February 2: GOES-10 image frame breaks due to interference.

At 1551z on February 3: GOES-8 image frame break due to address phasing

At 1635z and 1640z on February 4: GOES-8 image frame breaks due to ground
system processor problems.




Questions or comments to

SpaceRef staff editor.